[figure: 9.7 cm]

SUPERFLUID CHARACTERISTICS OF THE TWO-DIMENSIONAL PENSON-KOLB MODEL

Wojciech R. Czart and Stanisław Robaszkiewicz Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland

We study superconducting properties of the two-dimensional (d = 2) Penson-Kolb model, i.e. the tight-binding model with the pair-hopping (intersite charge exchange) interaction J, and analyze the effects of phase fluctuations on the s-wave superconductivity in this system. The evolution of the critical fields, the coherence length, the Ginzburg ratio and the London penetration depth λ_{L} with particle concentration n and pairing strength is determined. Within the Kosterlitz-Thouless scenario the critical temperatures T_{c} are calculated and compared with the ones from BCS-HFA T_{HF A}. The results show the enhancement of the gap to critical temperature ratio due to the phase fluctuations as well as a separation of the energy scales for the pair formation (∼ k_{B}T_{HF A}) and the phase coherence (∼ k_{B}T_{C}). The Uemura type plots, i.e. the T_{c} vs λ_{L}^{−2}(0), are obtained both as a function of n and as a function of pairing strength and they are compared with those found earlier for the d = 2 attractive Hubbard model and the d = 2 extended Hubbard model with intersite attraction.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Wojciech R. Czart

Address for correspondence : Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań

Email address : czart@amu.edu.pl